OpenStudy (anonymous):

Annual profit in thousands of dollars is given by the function, P(x) = -.01x2 + 50x + 30,000, where x is the number of items sold, x ≥ 0. will this profit function have a maximum? if so what would it be?

OpenStudy (anonymous):

would i use the formula x=-b/2a

OpenStudy (anonymous):

you can derive it to get the maximum y'=-.02x+50 try getting the critical points -.02x + 50 = 0 50=.02x x= 2500 get the y y= -.01(2500)^2 + 50(2500)+30000 y= -62500 + 125000 + 30000 y= 92500 the function has a maximum 92500 at x = 2500

OpenStudy (anonymous):

that is what i got

OpenStudy (anonymous):

so the max weekly profit is approx 92,500 when 2,500 units are sold

OpenStudy (anonymous):

yes ^^

OpenStudy (anonymous):

so then what steps should the company do to prepare for my answer which was 92,500 was the max weekly profit when 2,500 units are sold?

OpenStudy (anonymous):

can someone help me with this one question?